### v1

# thanks: Corresponding authorthanks: Corresponding author

Abstract
One of the most fundamental relationships in modern cosmology is the cosmic distance duality relation (CDDR), which describes the relationship between the angular diameter distance ($D_{\rm A}$) and the luminosity distance ($D_{\rm L}$), and is expressed as: $\eta(z)=D_{\rm L}(z)(1+z)^{-2}/D_{\rm A}(z)=1$. In this work, we conduct a comprehensive test of the CDDR by combining baryon acoustic oscillation (BAO) data from the SDSS and DESI surveys with type Ia supernova (SN) data from PantheonPlus and DESY5. We utilize an artificial neural network approach to match the SN and BAO data at the same redshift. To explore potential violations of the CDDR, we consider three different parameterizations: (i) $\eta(z)=1+\eta_{0}z$; (ii) $\eta(z)=1+\eta_{0}z/(1+z)$; (iii) $\eta(z)=1+\eta_{0}\ln(1+z)$. Our results indicate that the calibration of the SN absolute magnitude $M_{\rm B}$ plays a crucial role in testing potential deviations from the CDDR, as there exists a significant negative correlation between $\eta_{0}$ and $M_{\rm B}$. For PantheonPlus analysis, when $M_{\rm B}$ is treated as a free parameter, no evidence of CDDR violation is found. In contrast, fixing $M_{\rm B}$ to the $M_{\rm B}^{\rm D20}$ prior with $-19.230\pm 0.040$ mag leads to a deviation at approximately the $2\sigma$ level, while fixing $M_{\rm B}$ to the $M_{\rm B}^{\rm B23}$ prior with $-19.396\pm 0.016$ mag remains in agreement with the CDDR. Furthermore, overall analyses based on the SDSS+DESY5 and DESI+DESY5 data consistently show no evidence of the deviation from the CDDR.
I Introduction
The cosmic distance duality relation (CDDR), also known as the Etherington relation Etherington (1933), is a fundamental relationship in modern cosmology. This relationship is predicated on three fundamental assumptions: the description of space-time by the metric theory of gravity, the photons propagate along null geodesics, and the conservation of the photon number Bassett and Kunz (2004). Specifically, for a given redshift, the relation can be expressed as $D_{\rm L}(z)=(1+z)^{2}D_{\rm A}(z)$, which establishes a connection between two critical measures of cosmic distance for an astronomical object: the luminosity distance ($D_{\rm L}$) and the angular diameter distance ($D_{\rm A}$). Therefore, testing the validity of the CDDR is essential for verifying the theories of gravity and the foundations of modern cosmology Avgoustidis et al. (2010); Gonçalves et al. (2012).
[…]
In recent years, a considerable amount of research has employed various astronomical observations to test the CDDR Uzan et al. (2004); De Bernardis et al. (2006); Li et al. (2011); Nair et al. (2011); Holanda et al. (2010); Liang et al. (2013); Meng et al. (2012); Khedekar and Chakraborti (2011); Goncalves et al. (2012); Lima et al. (2011); Cardone et al. (2012); Yang et al. (2013); Wu et al. (2015); Ma and Corasaniti (2018); Holanda et al. (2017); More et al. (2016); Rana et al. (2017); Li and Lin (2018); Lin et al. (2018); Qi et al. (2019); Holanda et al. (2019); Zhou et al. (2021); Qin et al. (2021); Bora and Desai (2021); Mukherjee and Mukherjee (2021); Liu et al. (2021); Xu et al. (2022); Tonghua et al. (2023); Yang et al. (2024); Tang et al. (2025); Jesus et al. (2025); Qi et al. (2025); Alfano and Luongo (2025). Type Ia supernova (SN) have been widely used to determine the luminosity distance, whereas the angular diameter distance, is typically obtained through multiple observational techniques, including (but not limited to) the Sunyaev–Zeldovich effect and gas mass fraction measurements in galaxy clusters Holanda et al. (2010); Li et al. (2011), strong gravitational lensing systems Liao et al. (2022); Ruan et al. (2018); Qi et al. (2025), the angular size of ultra-compact radio sources Li and Lin (2018), and baryon acoustic oscillations (BAO) Wu et al. (2015). In particular, SN and BAO observations play an important role in testing the CDDR Wu et al. (2015); Xu and Huang (2020); Wang et al. (2024); Xu et al. (2022). For example, Ref. Wu et al. (2015) tested the CDDR by comparing the SN data from Union 2.1 compilation with five BAO measurements, and concluded that the high precision of BAO measurements makes them an effective tool for validating the CDDR. Ref. Wang et al. (2024) tested the CDDR relation by utilizing SN data from the Pantheon sample and transverse BAO measurements from Sloan Digital Sky Survey (SDSS), demonstrating that transverse BAO measurements can serve as a powerful tool for validating the CDDR relation.
Recently, the Dark Energy Spectroscopic Instrument (DESI) publicly released its second data release (DR2), which includes measurements of BAO from over 14 million extragalactic sources, spanning the redshift range $0.295\leq z\leq 2.330$. This release has marked the beginning of a new era in high-precision BAO measurements. The DESI DR2 BAO data, combined with SN data from PantheonPlus, Union3, and DESY5, along with CMB data from Planck and ACT, provide more stringent constraints on the dark energy equation of state, revealing a $2.8\sigma-4.2\sigma$ preference for dynamical dark energy Abdul Karim et al. (2025). DESI BAO data have attracted a significant number of researchers seeking to constrain various aspects of cosmological physics, such as dark energy Li et al. (2024); Giarè et al. (2024); Dinda and Maartens (2025); Escamilla et al. (2024); Sabogal et al. (2024); Li et al. (2025a); Li and Wang (2024); Wang and Piao (2024); Giarè et al. (2024); Huang et al. (2025b); Li et al. (2025b); Wu (2025); Li and Zhang (2025); Li and Wang (2025); Barua and Desai (2025); Yashiki (2025); Ling et al. (2025); Goswami and Das (2025); Yang et al. (2025a); Pang et al. (2025); You et al. (2025); Özülker et al. (2025); Cheng et al. (2025); Pan et al. (2025); Qiang et al. (2025); Liu et al. (2025a); Gialamas et al. (2025), non-cold dark matter Yang et al. (2025b); Chen and Loeb (2025); Wang (2025); Kumar et al. (2025); Abedin et al. (2025a); Khoury et al. (2025); Araya et al. (2025); Li et al. (2025c, d); Paliathanasis (2025), neutrino Du et al. (2025a); Jiang et al. (2025); Du et al. (2025b); Feng et al. (2025); Roy Choudhury (2025), and modified gravity Cai et al. (2025); Li et al. (2025e); Odintsov et al. (2025); Nojiri et al. (2025). Additionally, it is essential to revisit the CDDR in the context of the most recent high-precision BAO data from DESI, as emphasized by recent studies Zhang et al. (2025b); Wang et al. (2025); Santos et al. (2025); Afroz and Mukherjee (2025); Yang et al. (2025c); Keil et al. (2025); Teixeira et al. (2025). For example, Favale et al. (2024) investigated the consistency of the CDDR by combining SDSS and DESI mixed BAO data with SN data, reporting indications of a potential deviation at approximately the $2\sigma$ level. The authors attributed this deviation to possible tensions between the SN and BAO data, or potentially internal inconsistencies within the BAO data subsets.
In this work, we test the CDDR using BAO data from DESI and SDSS, along with SN data from PantheonPlus and DESY5. We utilize the artificial neural network (ANN) approach to align the SN and BAO data at the same redshifts. Our aim is to utilize the pairwise independent combinations of the BAO and SN datasets to provide a comprehensive and robust analysis of the CDDR and investigate any potential violations of this fundamental relation.
This work is organized as follows. In Sec. II, we describe the BAO and SN data used in our analysis and outline the methodology for testing the CDDR. In Sec. III, we report the constraint results and make some relevant discussions. The conclusion is given in Sec. IV.
II methodology and data
We employ the function $\eta(z)$ to probe potential violations of the CDDR at various redshifts by comparing the luminosity distance from SN with the angular diameter distance derived from the BAO measurements. The definition of $\eta(z)$ is given by
$\eta(z)=\frac{D_{\rm L}}{D_{\rm A}(1+z)^{2}}.$ (1)
Any departure of $\eta(z)\neq 1$ at a given redshift signifies an inconsistency between the CDDR and the cosmological data. We parameterize $\eta(z)$ using three forms: (i) P1: $\eta(z)=1+\eta_{0}\,z$; (ii) P2: $\eta(z)=1+\eta_{0}\,z/(1+z)$; (iii) P3: $\eta(z)=1+\eta_{0}\,\ln(1+z)$.
The observed value $\eta_{\rm obs}(z)$ is computed using Eq. (1), and according to the error propagation formula, its uncertainty is given by
$\sigma^{2}_{\eta_{\rm obs}}=\eta_{\rm obs}^{2}\left[\biggl{(}\frac{\sigma_{D_{\rm A}(z)}}{D_{\rm A}(z)}\biggr{)}^{2}+\biggl{(}\frac{\sigma_{D_{\rm L}(z)}}{D_{\rm L}(z)}\biggr{)}^{2}\right].$ (2)
Finally, the best-fit for the parameter $\eta_{0}$ is evaluated through
$\chi^{2}(\eta_{0})=\sum_{i=1}^{N}\frac{\bigl{[}\eta(z_{i})-\eta_{\rm obs,\,i}(z_{i})\bigr{]}^{2}}{\sigma^{2}_{\eta_{\rm obs,\,i}}}\,.$ (3)
Table 1: The BAO samples utilized in this work including DESI DR2 data and SDSS data.
Tracer ${z_{\mathrm{eff}}}$ ${D_{\mathrm{M}}/r_{\mathrm{d}}}$ References
DESI
DR2 LRG1 0.51 13.59 $\pm$ 0.17 Abdul Karim et al. (2025)
DR2 LRG2 0.71 17.35 $\pm$ 0.18 Abdul Karim et al. (2025)
DR2 LRG3 + ELG1 0.93 21.58 $\pm$ 0.15 Abdul Karim et al. (2025)
DR2 ELG2 1.32 27.60 $\pm$ 0.32 Abdul Karim et al. (2025)
DR2 QSO 1.48 30.51 $\pm$ 0.76 Abdul Karim et al. (2025)
SDSS
BOSS LRG1 0.38 10.27 $\pm$ 0.15 Gil-Marín et al. (2017)
BOSS LRG2 0.51 13.38 $\pm$ 0.18 Gil-Marín et al. (2017)
eBOSS LRG 0.70 17.65 $\pm$ 0.30 Bautista et al. (2020)
eBOSS ELG 0.85 19.50 $\pm$ 1.00 Hou et al. (2020)
eBOSS QSO 1.48 30.21 $\pm$ 0.79 Hou et al. (2020)
The data of $D_{\rm L}$ we use in this work are obtained from following two SN datasets:
  • PantheonPlus: The PantheonPlus comprises 1550 spectroscopically confirmed supernovae from 18 different surveys, covering the redshift range $2.260.01<z<$ 2.26 Brout et al. (2022) [^2] Data available at https://github.com/PantheonPlusSH0ES/DataRelease.
  • DESY5: The DESY5 sample comprises 1635 photometrically classified supernovae from the released part of the full 5 yr data of the Dark Energy Survey collaboration (with the redshift range $1.30.1<z<$ 1.3), complemented by 194 low-redshift supernovae from the CfA3 Hicken et al. (2009), CfA4 Hicken et al. (2012), CSP Krisciunas et al. (2017), and Foundation Foley et al. (2018) samples (with redshifts in the range $0.10.025<z<$ 0.1), for a total of 1829 supernovae Abbott et al. (2024). [^3] Data available at https://github.com/des-science/DES-SN5YR.
The SN distance modulus $\mu$ is given by $\mu=m_{\rm B}-M_{\rm B}$, where $m_{\rm B}$ represents the observed peak apparent magnitude in the rest-frame $B$ band, $M_{\rm B}$ is the absolute magnitude. The luminosity distance $D_{\rm L}(z)$ connects to $\mu(z)$ through
$\mu(z)=5\log_{10}\biggl{[}\frac{D_{\rm L}(z)}{\mathrm{Mpc}}\biggr{]}+25,$ (4)
with the corresponding uncertainty in $D_{\rm L}$ derived by standard error propagation
$\sigma_{D_{\rm L}}=\frac{D_{\rm L}\ln 10}{5}\sigma_{\mu}.$ (5)
Figure 1: The ANN reconstruction for the SN data. Left panel: The ANN-reconstructed $M_{\rm B}$ for the PantheonPlus over the redshift range $0.01\leq z\leq 2.26$. The light green shaded band denotes the $1\sigma$ confidence interval of the ANN reconstruction, while the black points with error bars correspond to the PantheonPlus data points. Right panel: The ANN-reconstructed $\mu$ for the DESY5 over the redshift range $0.025\leq z\leq 1.3$. The light red shaded band denotes the $1\sigma$ confidence interval of the ANN reconstruction, and the light grey points with error bars represent the DESY5 data points.
The distance measurements of SN crucially rely on the calibration of $M_{\rm B}$. For the PantheonPlus analysis, we firstly perform fits treating $M_{\rm B}$ as a free parameter. Additionally, to evaluate the impacts of different prior values of $M_{\rm B}$ on the CDDR test, we adopt two specific Gaussian priors: (i) $M_{\rm B}^{\rm D20}=-19.230\pm 0.040\,\mathrm{mag}$, from low‑$z$ SN ($0.150.023<z<$ 0.15) in $\Lambda$CDM via de‑marginalization of the SH0ES result Camarena and Marra (2020); Reid et al. (2019); (ii) $M_{\rm B}^{\rm B23}=-19.396\pm 0.016\,\mathrm{mag}$, from combined SN and BAO data Dinda and Banerjee (2023). Including the observational error in $M_{\rm B}$, the uncertainty in the distance modulus is $\sigma_{\mu}=\sqrt{\sigma_{M_{\rm B}}^{2}+\sigma_{m_{\rm B}}^{2}}\,$. Conversely, the DESY5 compilation provides $\mu$ and $\sigma_{\mu}$ directly, as $M_{\rm B}$ is already normalized within the calibration. Therefore, there is no prior on $M_{\rm B}$ is imposed for DESY5 analysis in this work.
For the measurement of $D_{\rm A}(z)$, we utilize the transverse comoving distance $D_{\mathrm{M}}(z)/r_{\mathrm{d}}$ derived from the BAO observations of the DESI DR2 Abdul Karim et al. (2025) and SDSS Gil-Marín et al. (2017); Hou et al. (2020); Bautista et al. (2020). The specific BAO measurements are listed in Table 1. The angular diameter distance $D_{\rm A}(z)$ is obtained through
$D_{\rm A}(z)=\frac{D_{\rm M}(z)}{1+z}.$ (6)
In this work, we adopt the sound horizon $r_{\mathrm{d}}=147.78~{}\mathrm{Mpc}$ corresponding to the Planck 2018 result Aghanim et al. (2020). It is also worth noting that, due to the lower redshift range of the DESY5 data, we do not include the DESI DR2 ELG2, QSO measurements, and the SDSS eBOSS QSO data point in the DESY5-related analysis.
Table 2: Fitting results ($1\sigma$ confidence level) in the $P_{1}$, $P_{2}$, and $P_{3}$ models from the BAO and SN data.
Data $P_{1}$ model $P_{2}$ model $P_{3}$ model
$\eta_{0}$ $M_{\rm B}$ $\eta_{0}$ $M_{\rm B}$ $\eta_{0}$ $M_{\rm B}$
SDSS + PantheonPlus $0.040^{+0.130}_{-0.190}$ $-19.510^{+0.290}_{-0.230}$ $0.290^{+0.360}_{-0.810}$ $-19.630^{+0.580}_{-0.400}$ $0.100^{+0.220}_{-0.390}$ $-19.540^{+0.400}_{-0.300}$
DESI + PantheonPlus $0.060^{+0.120}_{-0.210}$ $-19.560^{+0.420}_{-0.290}$ $0.480^{+0.440}_{-0.990}$ $-19.810^{+0.790}_{-0.570}$ $0.170^{+0.220}_{-0.480}$ $-19.650^{+0.570}_{-0.390}$
SDSS + PantheonPlus ($M_{\rm B}^{\rm D20}$) $-0.120\pm 0.067$ $-19.230\pm 0.040$ $-0.240\pm 0.130$ $-19.230\pm 0.040$ $-0.178\pm 0.093$ $-19.230\pm 0.040$
DESI + PantheonPlus ($M_{\rm B}^{\rm D20}$) $-0.098\pm 0.052$ $-19.230\pm 0.040$ $-0.220\pm 0.113$ $-19.230\pm 0.040$ $-0.162\pm 0.068$ $-19.230\pm 0.040$
SDSS + PantheonPlus ($M_{\rm B}^{\rm B23}$) $-0.040\pm 0.071$ $-19.396\pm 0.016$ $-0.052\pm 0.131$ $-19.396\pm 0.016$ $-0.050\pm 0.102$ $-19.396\pm 0.016$
DESI + PantheonPlus ($M_{\rm B}^{\rm B23}$) $-0.035\pm 0.056$ $-19.396\pm 0.016$ $-0.070\pm 0.124$ $-19.396\pm 0.016$ $-0.043\pm 0.088$ $-19.396\pm 0.016$
SDSS + DESY5 $-0.035\pm 0.074$- $-0.080\pm 0.120$- $-0.059\pm 0.093$-
DESI + DESY5 $-0.052\pm 0.066$- $-0.090\pm 0.110$- $-0.068\pm 0.085$-
Figure 2: The triangular plots on $\eta_{0}$ and $M_{\rm B}$ for the $P_{1}$ (left panel), $P_{2}$ (middle panel), and $P_{3}$ (right panel) models, using SDSS, DESI, and PantheonPlus data.
In order to match the BAO and SN data at a given redshift, we reconstruct $m_{\rm B}$ or $\mu$ of SN using a nonparametric ANN method implemented with the REFANN Python package (Wang et al., 2020). The ANN comprises a single hidden layer of 4096 neurons and is trained on the PantheonPlus and DESY5 datasets to yield a data‑driven reconstruction of $m_{\rm B}$ or $\mu$. From this reconstruction we extract $D_{\rm L}(z)$ at the redshifts of BAO measurements and compare it with the angular diameter distance $D_{\rm A}(z)$ inferred from the same BAO data to test the CDDR. The ANN reconstruction $m_{\rm B}$ for PantheonPlus and $\mu$ for DESY5 in this work are shown in Fig. 1.
This nonparametric framework does not assume any predefined functional form, providing a flexible approach for probing potential deviations from the standard cosmological model Wang et al. (2021); Liu et al. (2021); Qi et al. (2023); Liu et al. (2025b); Abedin et al. (2025b). The precision of the reconstructed $m_{\rm B}$ or $\mu$ and the width of their confidence intervals depend on hyperparameter choices, such as network architecture and training strategy, which influence the expressive power and generalization capability of ANN. Nonetheless, the ANN‑based reconstruction remains a powerful, model‑independent tool for evaluating the CDDR using BAO and SN data.
We perform Bayesian analysis using the publicly available package Cobaya [^4] https://github.com/CobayaSampler/cobaya Torrado and Lewis (2021) and assess the convergence of the Markov Chain Monte Carlo (MCMC) Lewis and Bridle (2002); Lewis (2013) chains using the Gelman-Rubin statistics quantity $R-1<0.01$ Gelman and Rubin (1992). The MCMC chains are analyzed using GetDist [^5] https://github.com/cmbant/getdist/ Lewis (2019).
III Results and discussions
Figure 3: The 1D marginalized posterior distributions of the $\eta_{0}$ parameter for the $P_{1}$, $P_{2}$, and $P_{3}$ models of the CDDR with $M_{\rm B}$ fixed, using $M_{\rm B}^{\rm B23}$ (upper panel) and $M_{\rm B}^{\rm D20}$ (lower panel), based on SDSS, DESI, and PantheonPlus data.
In this section, we present the constraints on $\eta_{0}$ and $M_{\rm B}$ for the $P_{1}$, $P_{2}$, and $P_{3}$ parameterizations of $\eta(z)$, using the SDSS, DESI, PantheonPlus, and DESY5 data. The marginalized results with $1\sigma$ errors are summarized in Table 2. The triangular plots, along with the one-dimensional (1D) marginalized distributions of $\eta_{0}$ and $M_{\rm B}$ based on various datasets and parameterizations, are shown in Figs. 24.
In Fig. 2, we show the triangular plot of the constraint results obtained by treating $M_{\rm B}$ as a free parameter, using the SDSS, DESI, and PantheonPlus data. The constraints on $\eta_{0}$ are $0.060^{+0.120}_{-0.210}$ ($P_{1}$ model), $0.480^{+0.440}_{-0.990}$ ($P_{2}$ model), and $0.170^{+0.220}_{-0.480}$ ($P_{3}$ model), based on the DESI+PantheonPlus data. This indicates that in all parameterized models, there is no evidence for a violation of the CDDR. This result emphasizes the robustness of our CDDR analysis, and the consistency among the three parameterized models $P_{1}$, $P_{2}$, and $P_{3}$ also suggests that using multiple parameterizations of $\eta(z)$ can provide complementary insights into potential deviations from the CDDR. Similarly, the SDSS+PantheonPlus combination yields consistent results, indicating good agreement between the BAO measurements from SDSS and DESI DR2. The constraints on $M_{\rm B}$ using the DESI+PantheonPlus datasets are $-19.560^{+0.420}_{-0.290}$ ($P_{1}$ model), $-19.810^{+0.790}_{-0.570}$ ($P_{2}$ model), and $-19.650^{+0.570}_{-0.390}$ ($P_{3}$ model). The central values of $M_{\rm B}$ in all parameterized models are slightly lower than the commonly accepted value of approximately $-19.3$ mag. This slight shift may result from the inclusion of the possible deviation parameter $\eta_{0}$ in the CDDR analysis.
Figure 4: The 1D marginalized posterior distributions of the $\eta_{0}$ parameter for the $P_{1}$, $P_{2}$, and $P_{3}$ models of the CDDR, using SDSS, DESI, and DESY5 data.
It is worth noting that there is a strong negative correlation between $\eta_{0}$ and $M_{\rm B}$, as shown in Fig. 2. This correlation indicates that, different $M_{\rm B}$ priors will have a significant impact on our detection of potential deviations from the CDDR. Therefore, we further adopt the $M_{\rm B}^{\rm D20}$ and $M_{\rm B}^{\rm B23}$ priors to assess the potential impact of $M_{\rm B}$ on deviations from the CDDR. The 1D marginalized posterior distributions of $\eta_{0}$ for $P_{1}$, $P_{2}$, and $P_{3}$ models are shown in Fig. 3. When adopting the $M_{\rm B}^{\rm D20}$ prior with $M_{\rm B}=-19.230\pm 0.040$ mag, $P_{1}$, $P_{2}$, and $P_{3}$ models yield $\eta_{0}=-0.098\pm 0.052$, $\eta_{0}=-0.220\pm 0.113$, and $\eta_{0}=-0.162\pm 0.068$ using DESI+PantheonPlus data, respectively. All three models consistently show a deviation from CDDR at approximately $2\sigma$ confidence level, especially for the $P_{3}$ model, which has reached $2.4\sigma$. In contrast, if we adopt $M_{\rm B}^{\rm B23}$ prior with $M_{\rm B}=-19.396\pm 0.016$ mag, the constraint values of $\eta_{0}$ are $-0.035\pm 0.056$ ($P_{1}$ model), $-0.070\pm 0.124$ ($P_{2}$ model), and $-0.043\pm 0.088$ ($P_{3}$ model), suggesting that no deviation from the CDDR is found in all models. This indicates that if a fixed $M_{\rm B}$ prior is adopted, a bias of approximately 0.2 mag of $M_{\rm B}$ can lead to a deviation of $2\sigma$ in CDDR, emphasizing the necessity of analysing the impact of $M_{\rm B}$ prior on CDDR deviation. Furthermore, we can obtain that the absolute magnitude of the SN needs to be calibrated to approximately $-19.4$ mag in order for the CDDR to remain undeviation using the DESI and PantheonPlus data.
Next, we present the 1D marginalized posterior distributions of $\eta_{0}$ for the $P_{1}$, $P_{2}$, and $P_{3}$ models, obtained using the DESI+DESY5 and SDSS+DESY5 data, as shown in Fig. 4. Using the SDSS+DESY5 data, the constraint values of $\eta_{0}$ are $-0.035\pm 0.074$, $-0.080\pm 0.120$, and $-0.059\pm 0.093$ for $P_{1}$, $P_{2}$, and $P_{3}$ models, respectively. Similarly, the constraint values of $\eta_{0}$ are $-0.052\pm 0.066$ ($P_{1}$ model), $-0.090\pm 0.110$ ($P_{2}$ model), and $-0.068\pm 0.085$ ($P_{3}$ model) using DESI+DESY5 data, respectively. We find that these results consistently show no deviation from the CDDR in all parameterized models. Therefore, except for the case of adopting $M_{\rm B}^{\rm D20}$ prior, we did not detect any deviation from CDDR in other cases, indicating that the current data is well consistent with CDDR.
Finally, it is necessary to compare the constraints on $\eta_{0}$ given by different observations to explore the capabilities of BAO and SN data. The constraints of $\eta_{0}$ using DESI+DESY5 are approximately 36%, 56%, and 47% higher than that obtained using the SDSS eBOSS DR16 BAO dataset combined with the Pantheon SN sample, for $P_{1}$, $P_{2}$, and $P_{3}$ models, respectively Xu et al. (2022). Similarly, the constraints are approximately 49%, 71%, 72% more restrictive than those obtained from six well-studied strong gravitational lensing systems and PantheonPlus SN data for three models Qi et al. (2025). Furthermore, this results are comparable to those inferred from compact radio quasars measurements and Pantheon SN samples Yang et al. (2024). Therefore, the latest BAO and SN data provide a more effective and accurate way to test CDDR.
IV Conclusion
The CDDR is a fundamental relationship of modern cosmology, and its validation largely relies on precise distance measurements. In this work, we explore the potential deviation of the CDDR using three phenomenological parameterizations of the function $\eta(z)$, by combining BAO data from SDSS and DESI DR2 with SN data from PantheonPlus and DESY5. We also evaluate the impacts of the SN absolute magnitude on testing the CDDR.
When $M_{\rm B}$ is treated as a free parameter, we find a strong negative correlation between $\eta_{0}$ and $M_{\rm B}$, indicating that the uncertainty in $M_{\rm B}$ directly affects the inferred CDDR deviation. Meanwhile, all three parameterizations show no evidence for a deviation from CDDR using the SDSS, DESI, and PantheonPlus data. The best-fit values of $M_{\rm B}$ are slightly lower than the commonly adopted value of $-19.3$ mag, suggesting that the inclusion of a deviation parameter in CDDR can lead to a small shift in $M_{\rm B}$ from its conventional calibration. When fixing the $M_{\rm B}^{\rm D20}$ prior ($-19.230\pm 0.040$ mag), we find a deviation from CDDR at the approximately $2\sigma$ level, especially in the $P_{3}$ model, where the deviation reaches $2.4\sigma$ level. In contrast, the $M_{\rm B}^{\rm B23}$ prior ($-19.396\pm 0.016$ mag) shows no deviation from CDDR in all models. This implies that a shift of approximately 0.2 mag in $M_{\rm B}$ can lead to a $2\sigma$ level deviation from CDDR. Finally, the joint analysis of DESI, SDSS, and DESY5 data maintains CDDR across all models.
In conclusion, our analysis has conducted a meticulous examination of the CDDR, revealing that the absolute magnitude $M_{\rm B}$ of SN has a significant impact on any deviation from the CDDR. Therefore, it is crucial in future to employ more data to calibrate $M_{\rm B}$ with higher accuracy and precision, thereby determining its value and re‐evaluating the CDDR.
Acknowledgments
We thank Tian-Yang Sun, Sheng-Han Zhou, Jia-Le Ling, Yi-Min Zhang, and Si-Ren Xiao for their helpful discussions. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12473001, 11975072, 11875102, 11835009, and 12205039), the China Manned Space Program (Grant No. CMS-CSST-2025-A02), and the National 111 Project (Grant No. B16009).