### v2

# thanks: Corresponding authorthanks: Corresponding author

Abstract
One of the most fundamental relationships in modern cosmology is the cosmic distance duality relation (CDDR), which describes the relationship between the angular diameter distance ($D_{\rm A}$) and the luminosity distance ($D_{\rm L}$), and is expressed as $\eta(z)=D_{\rm L}(z)(1+z)^{-2}/D_{\rm A}(z)=1$. In this work, we conduct a comprehensive test of the CDDR by combining baryon acoustic oscillation (BAO) data from the SDSS and DESI surveys with type Ia supernova (SN) data from PantheonPlus and DESY5. We utilize an artificial neural network approach to match the SN and BAO data at the same redshift. To explore potential violations of the CDDR, we consider three different parameterizations: (i) $\eta(z)=1+\eta_{0}z$; (ii) $\eta(z)=1+\eta_{0}z/(1+z)$; (iii) $\eta(z)=1+\eta_{0}\ln(1+z)$. Our results indicate that the calibration of the SN absolute magnitude $M_{\rm B}$ plays a crucial role in testing potential deviations from the CDDR, as there exists a significant negative correlation between $\eta_{0}$ and $M_{\rm B}$. For PantheonPlus analysis, when $M_{\rm B}$ is treated as a free parameter, no evidence of CDDR violation is found. In contrast, fixing $M_{\rm B}$ to the $M_{\rm B}^{\rm D20}$ prior with $19.230\pm 0.040$ mag leads to a deviation at approximately the $2\sigma$ level, while fixing $M_{\rm B}$ to the $M_{\rm B}^{\rm B23}$ prior with $19.396\pm 0.016$ mag remains in agreement with the CDDR. Furthermore, overall analyses based on the SDSS+DESY5 and DESI+DESY5 data consistently show no evidence of deviation from the CDDR.
I Introduction
The cosmic distance duality relation (CDDR), also known as the Etherington relation Etherington1933, is a fundamental relationship in modern cosmology. This relationship is predicated on three fundamental assumptions: the description of spacetime by the metric theory of gravity, the propagation of photons along null geodesics, and the conservation of the photon number Bassett:2003vu. Specifically, for a given redshift, the relation can be expressed as $D_{\rm L}(z)=(1+z)^{2}D_{\rm A}(z)$, which establishes a connection between two critical measures of cosmic distance for an astronomical object: the luminosity distance ($D_{\rm L}$) and the angular diameter distance ($D_{\rm A}$). Therefore, testing the validity of the CDDR is essential for verifying the theories of gravity and the foundations of modern cosmology Avgoustidis2010; Holanda2011.
[…]
In recent years, a considerable amount of research has employed various astronomical observations to test the CDDR Uzan:2004my; DeBernardis:2006ii; Li:2011exa; Nair:2011dp; Holanda:2010vb; Liang:2011gm; Meng:2011nt; Khedekar:2011gf; Goncalves:2011ha; Lima:2011ye; Cardone:2012vd; Yang:2013coa; Wu:2015ixa; Ma:2016bjt; Holanda:2016msr; More:2016fca; Rana:2017sfr; Li:2017zrx; Lin:2018qal; Qi:2019spg; Holanda:2019vmh; Zhou:2020moc; Qin:2021jqy; Bora:2021cjl; Mukherjee:2021kcu; Liu:2021fka; Xu:2022zlm; Tonghua:2023hdz; Yang:2024icv; Tang:2024zkc; Jesus:2024nrl; Qi:2024acx; Alfano:2025gie. Type Ia supernova (SN) have been widely used to determine the luminosity distance, whereas the angular diameter distance, is typically obtained through multiple observational techniques, including (but not limited to) the Sunyaev–Zeldovich effect and gas mass fraction measurements in galaxy clusters Holanda:2010vb; Li:2011exa, strong gravitational lensing systems Liao:2022gde; Ruan:2018dls; Qi:2024acx, the angular size of ultra-compact radio sources Li:2017zrx, and baryon acoustic oscillations (BAO) Wu:2015ixa. In particular, SN and BAO observations play an important role in testing the CDDR Wu:2015ixa; Xu:2020fxj; Wang:2024rxm; Xu:2022zlm. For example, Ref. Wu:2015ixa tested the CDDR by comparing the SN data from the SCP Union2.1 compilation with five BAO measurements, and concluded that the high precision of BAO measurements makes them an effective tool for validating the CDDR. Reference Wang:2024rxm tested the CDDR relation by utilizing SN data from the Pantheon sample and transverse BAO measurements from the Sloan Digital Sky Survey (SDSS), demonstrating that transverse BAO measurements can serve as a powerful tool for validating the CDDR relation.
Recently, the Dark Energy Spectroscopic Instrument (DESI) publicly released its second data release (DR2), which includes measurements of BAO from over 14 million extragalactic sources, spanning the redshift range $0.295\leq z\leq 2.330$. This release has marked the beginning of a new era in high-precision BAO measurements. The DESI DR2 BAO data, combined with SN data from PantheonPlus, Union3, and DESY5, along with CMB data from Planck and ACT, provide more stringent constraints on the dark energy equation of state, revealing a $2.8\sigma-4.2\sigma$ preference for dynamical dark energy DESI:2025zgx. DESI BAO data have attracted a significant number of researchers seeking to constrain various aspects of cosmological physics, such as dark energy Li:2024qso; Giare:2024gpk; Dinda:2024ktd; Colgain:2024xqj; Escamilla:2024ahl; Colgain:2024ksa; Sabogal:2024yha; Colgain:2024mtg; Li:2024qus; Li:2024bwr; Wang:2024dka; Giare:2024smz; Huang:2025som; Li:2025owk; Wu:2025wyk; Li:2025ula; Li:2025ops; Barua:2025ypw; Yashiki:2025loj; Ling:2025lmw; Goswami:2025uih; Yang:2025boq; Colgain:2025nzf; Pang:2025lvh; You:2025uon; Ozulker:2025ehg; Cheng:2025lod; Pan:2025qwy; Qiang:2025cxp; Liu:2025myr; Gialamas:2025pwv; Colgain:2025fct; Wu:2025vfs, non-cold dark matter Yang:2025ume; Chen:2025wwn; Wang:2025zri; Kumar:2025etf; Abedin:2025dis; Khoury:2025txd; Araya:2025rqz; Li:2025eqh; Li:2025dwz; Paliathanasis:2025xxm, neutrinos Du:2024pai; Jiang:2024viw; Du:2025iow; Feng:2025mlo; RoyChoudhury:2025dhe; Du:2025xes; Zhou:2025nkb, and modified gravity Cai:2025mas; Li:2025cxn; Odintsov:2025jfq; Nojiri:2025low. Additionally, it is essential to revisit the CDDR in the context of the most recent high-precision BAO data from DESI, as emphasized by several studies based on the first data release of DESI Favale:2024sdq; Yang:2025qdg; Keil:2025ysb; Teixeira:2025czm and more recent analyses using DESI DR2 data Zhang:2025qbs; Wang:2025gus; Santos:2025gjf; Afroz:2025iwo. For example, Ref. Favale:2024sdq investigated the consistency of the CDDR by combining SDSS and DESI mixed BAO data with SN data, reporting indications of a potential deviation at approximately the $2\sigma$ level. The authors attributed this deviation to possible tensions between the SN and BAO data, or potentially internal inconsistencies within the BAO data subsets.
However, a major challenge in testing the CDDR using the latest DESI DR2 data is the difficulty of obtaining matched angular diameter distance and luminosity distance at identical redshifts, as most observational datasets provide these quantities at different redshift points. To address this issue, several methods for reconstructing the distance have been proposed, such as linear and polynomial fitting Holanda:2015zpz and Gaussian process Hees:2014lfa. Specifically, Ref. Wang:2019vxv proposed a non-parametric approach based on an artificial neural network (ANN) to reconstruct functions directly from observational data. In recent years, several studies have applied the ANN method to test the CDDR; see, e.g., Refs. Liu:2021fka; Xu:2022zlm; Tang:2022ykd; Qi:2024acx; Yang:2025qdg.
In this work, our primary motivation is to test the validity of the CDDR by utilizing the latest DESI DR2 BAO and DESY5 SN data and by applying an ANN approach to align the BAO and SN data at the same redshifts. To obtain more comprehensive and robust results, we also incorporate the SDSS BAO and PantheonPlus SN data, and perform cross-checks using four pairwise independent combinations of the BAO and SN datasets.
This work is organized as follows. In Sect. II, we describe the BAO and SN data used in our analysis and outline the methodology for testing the CDDR. In Sect. III, we report the constraint results and make some relevant discussions. The conclusion is given in Sect. IV.
II Methodology and data
We employ the function $\eta(z)$ to probe potential violations of the CDDR at various redshifts by comparing the luminosity distance from SN with the angular diameter distance derived from the BAO measurements. The definition of $\eta(z)$ is given by
$\eta(z)=\frac{D_{\rm L}}{D_{\rm A}(1+z)^{2}}.$ (1)
Any departure of $\eta(z)\neq 1$ at a given redshift signifies an inconsistency between the CDDR and the cosmological data. We parameterize $\eta(z)$ using three forms: (i) $P_{1}$: $\eta(z)=1+\eta_{0}\,z$; (ii) $P_{2}$: $\eta(z)=1+\eta_{0}\,z/(1+z)$; (iii) $P_{3}$: $\eta(z)=1+\eta_{0}\,\ln(1+z)$. The first function represents the first-order term of a Taylor series expansion, but it exhibits undesirable behavior at high redshift. In contrast, the second function performs well even at high redshift, evolving more gradually than the first. The third function is expressed in a more complex logarithmic form. The reason for considering these three typical parameterizations is that they are widely used in many studies (see, e.g., Refs. Liu:2021fka; Xu:2022zlm; Qi:2024acx; Yang:2024icv), facilitating a comparison with our results. Furthermore, it is important to highlight that the motivation for considering multiple parameterizations stems from previous studies, which show that in any expansion of the form $\eta(z)=a+bf(z)$, the associated uncertainties are highly sensitive to the choice of the function $f(z)$ (see, e.g., Ref. Colgain:2021pmf). This inherent arbitrariness in a single parameterization is best addressed by considering multiple parameterizations, thus enabling a more robust and unbiased analysis.
The observed value $\eta_{\rm obs}(z)$ is computed using Eq. (1), and according to the error propagation formula, its uncertainty is given by
$\sigma^{2}_{\eta_{\rm obs}}=\eta_{\rm obs}^{2}\left[\biggl(\frac{\sigma_{D_{\rm A}(z)}}{D_{\rm A}(z)}\biggr)^{2}+\biggl(\frac{\sigma_{D_{\rm L}(z)}}{D_{\rm L}(z)}\biggr)^{2}\right].$ (2)
Finally, the best-fit for the parameter $\eta_{0}$ is evaluated through
$\chi^{2}(\eta_{0})=\sum_{i=1}^{N}\frac{\bigl[\eta(z_{i})-\eta_{\rm obs,\,i}(z_{i})\bigr]^{2}}{\sigma^{2}_{\eta_{\rm obs,\,i}}}\,.$ (3)
Table 1: The BAO samples utilized in this work including DESI DR2 data and SDSS data.
Tracer $z_{\mathrm{eff}}$} $D_{\mathrm{M}}/r_{\mathrm{d}}$} References
DESI
DR2 LRG1 0.51 13.59 $\pm$ 0.17 DESI:2025zgx
DR2 LRG2 0.71 17.35 $\pm$ 0.18 DESI:2025zgx
DR2 LRG3 + ELG1 0.93 21.58 $\pm$ 0.15 DESI:2025zgx
DR2 ELG2 1.32 27.60 $\pm$ 0.32 DESI:2025zgx
DR2 QSO 1.48 30.51 $\pm$ 0.76 DESI:2025zgx
SDSS
BOSS LRG1 0.38 10.27 $\pm$ 0.15 Gil-Marin:2016wya
BOSS LRG2 0.51 13.38 $\pm$ 0.18 Gil-Marin:2016wya
eBOSS LRG 0.70 17.65 $\pm$ 0.30 eBOSS:2020lta
eBOSS ELG 0.85 19.50 $\pm$ 1.00 eBOSS:2020gbb
eBOSS QSO 1.48 30.21 $\pm$ 0.79 eBOSS:2020gbb
The data of $D_{\rm L}$ we use in this work are obtained from following two SN datasets:
  • PantheonPlus: The PantheonPlus comprises 1550 spectroscopically confirmed supernovae from 18 different surveys, covering the redshift range $0.01<z<2.26$ Brout:2022vxf. [^2] https://github.com/PantheonPlusSH0ES/DataRelease.
  • DESY5: The DESY5 sample comprises 1635 photometrically classified supernovae from the released part of the full 5-year data of the Dark Energy Survey collaboration (with a redshift range $0.1<z<1.3$), complemented by 194 low-redshift supernovae from the CfA3 Hicken:2009df, CfA4 Hicken:2012zr, CSP Krisciunas:2017yoe, and Foundation Foley:2017zdq samples (with redshifts in the range $0.025<z<0.1$), for a total of 1829 supernovae DES:2024jxu. [^3] https://github.com/des-science/DES-SN5YR.
The SN distance modulus $\mu$ is given by $\mu=m_{\rm B}-M_{\rm B}$, where $m_{\rm B}$ represents the observed peak apparent magnitude in the rest-frame $B$ band, $M_{\rm B}$ is the absolute magnitude. The luminosity distance $D_{\rm L}(z)$ connects to $\mu(z)$ through
$\mu(z)=5\log_{10}\biggl[\frac{D_{\rm L}(z)}{\mathrm{Mpc}}\biggr]+25,$ (4)
with the corresponding uncertainty in $D_{\rm L}$ derived by standard error propagation
$\sigma_{D_{\rm L}}=\frac{D_{\rm L}\ln 10}{5}\sigma_{\mu}.$ (5)
Figure 1: The ANN reconstruction for the SN data. Left panel: The ANN-reconstructed $m_{\rm B}$ for the PantheonPlus over the redshift range $0.01\leq z\leq 2.26$. The light green shaded band denotes the $1\sigma$ confidence interval of the ANN reconstruction, while the black points with error bars correspond to the PantheonPlus data points. Right panel: The ANN-reconstructed $\mu$ for the DESY5 over the redshift range $0.025\leq z\leq 1.3$. The light red shaded band denotes the $1\sigma$ confidence interval of the ANN reconstruction, and the light grey points with error bars represent the DESY5 data points.
[…]
For the measurement of $D_{\rm A}(z)$, we utilize the transverse comoving distance $D_{\mathrm{M}}(z)/r_{\mathrm{d}}$ derived from the BAO observations of the DESI DR2 DESI:2025zgx and SDSS Gil-Marin:2016wya; eBOSS:2020gbb; eBOSS:2020lta. The specific BAO measurements are listed in Table 1. The angular diameter distance $D_{\rm A}(z)$ is obtained through
$D_{\rm A}(z)=\frac{D_{\rm M}(z)}{1+z}.$ (6)
In this work, we adopt the sound horizon $r_{\mathrm{d}}=147.78~\mathrm{Mpc}$ corresponding to the Planck 2018 result Planck:2018vyg. Note that we do not consider the uncertainty in $r_{\rm d}$, since it is much smaller than the uncertainties in $D_{\rm L}(z)$ and $D_{\rm A}(z)$ and is therefore negligible. Additionally, some DESI and SDSS BAO measurements are excluded from this analysis because their redshifts lie beyond the range probed by the SN samples. Specifically, when using the PantheonPlus SN sample, we omit BAO measurements from the Lyman-$\alpha$ forest in DESI DR2 and SDSS eBOSS. When using the DESY5 SN sample, we omit the DESI DR2 ELG2, QSO, and Lyman-$\alpha$ forest measurements, as well as the SDSS eBOSS QSO and Lyman-$\alpha$ forest data points.
Table 2: Fitting results ($1\sigma$ confidence level) in the $P_{1}$, $P_{2}$, and $P_{3}$ models from the BAO and SN data.
Data $P_{1}$ model $P_{2}$ model $P_{3}$ model
$\eta_{0}$ $M_{\rm B}$ $\eta_{0}$ $M_{\rm B}$ $\eta_{0}$ $M_{\rm B}$
SDSS + PantheonPlus $0.040^{+0.130}_{-0.190}$ $19.510^{+0.290}_{-0.230}$ $0.290^{+0.360}_{-0.810}$ $19.630^{+0.580}_{-0.400}$ $0.100^{+0.220}_{-0.390}$ $19.540^{+0.400}_{-0.300}$
DESI + PantheonPlus $0.060^{+0.120}_{-0.210}$ $19.560^{+0.420}_{-0.290}$ $0.480^{+0.440}_{-0.990}$ $19.810^{+0.790}_{-0.570}$ $0.170^{+0.220}_{-0.480}$ $19.650^{+0.570}_{-0.390}$
SDSS + PantheonPlus $\times M_{\rm B}^{\rm D20}$ $0.120\pm 0.067$ $19.230\pm 0.040$ $0.240\pm 0.130$ $19.230\pm 0.040$ $0.178\pm 0.093$ $19.230\pm 0.040$
DESI + PantheonPlus $\times M_{\rm B}^{\rm D20}$ $0.098\pm 0.052$ $19.230\pm 0.040$ $0.220\pm 0.113$ $19.230\pm 0.040$ $0.162\pm 0.068$ $19.230\pm 0.040$
SDSS + PantheonPlus $\times M_{\rm B}^{\rm B23}$ $0.040\pm 0.071$ $19.396\pm 0.016$ $0.052\pm 0.131$ $19.396\pm 0.016$ $0.050\pm 0.102$ $19.396\pm 0.016$
DESI + PantheonPlus $\times M_{\rm B}^{\rm B23}$ $0.035\pm 0.056$ $19.396\pm 0.016$ $0.070\pm 0.124$ $19.396\pm 0.016$ $0.043\pm 0.088$ $19.396\pm 0.016$
SDSS + DESY5 $035\pm 0.074 - 0.0$$80\pm 0.120 - 0.0$$059\pm 0.093 -$
DESI + DESY5 $052\pm 0.066 - 0.0$$90\pm 0.110 - 0.0$$068\pm 0.085 -$
In order to match the BAO and SN data at a given redshift, we reconstruct $m_{\rm B}$ or $\mu$ of SN using a nonparametric ANN method implemented with the REFANN Python package Wang:2019vxv. This ANN method is fully data-driven, allowing us to reconstruct a function from any type of data without assuming a specific parameterization of the function. An ANN typically consists of three main layers: the input layer, the hidden layer, and the output layer. The input layer consists of the features of the dataset, with each neuron representing a feature or variable. Since our goal is to reconstruct the functional relationship from the real data, we use all available data as the training input, without retaining a separate testing set, which is consistent with many similar studies (see, e.g., Refs. Qi:2024acx; Yang:2024icv). The neurons in the hidden layer receive outputs from the previous layer and transmit them to the neurons in the next layer. The hidden layers in an ANN are responsible for learning and data processing, and can consist of one or more layers. In our analysis, a single hidden layer connects the supernova redshift with the corresponding $m_{\rm B}$ or $\mu$ values, with a total of 4096 neurons in the hidden layer. This choice of the number of neurons is consistent with our previous work, providing a reasonable balance between accuracy and computational efficiency Qi:2023oxv; Qi:2024acx. The output layer is primarily responsible for generating the final result. We use a nonlinear activation function (exponential linear unit), with the remaining hyperparameters set to their default values as specified in Ref. Wang:2019vxv, and the mean absolute error is used as the loss function. The optimization method employed is gradient descent, which iteratively updates the loss value in the opposite direction of the current gradient to reduce the loss. The training process is performed over 30000 iterations to ensure the convergence of the loss function.
The ANN method is trained on the PantheonPlus and DESY5 datasets to yield a data-driven reconstruction of $m_{\rm B}$ or $\mu$. It is worth noting that for the SN data, we used only the diagonal elements of the covariance matrix to facilitate reconstruction using ANN. From this reconstruction we extract $D_{\rm L}(z)$ at the redshifts of BAO measurements and compare it with the angular diameter distance $D_{\rm A}(z)$ inferred from the same BAO data to test the CDDR. The ANN reconstruction $m_{\rm B}$ for PantheonPlus and $\mu$ for DESY5 in this work are shown in Fig. 1. Furthermore, we construct the SN Hubble diagram and compute the distance duality ratio $\eta(z)$, together with their uncertainties at the corresponding BAO redshifts using Eqs. (1), (2), and (5). The detailed results and discussion are presented in Appendix A.
This nonparametric framework does not assume any predefined functional form, providing a flexible approach for probing potential deviations from the standard cosmological model Wang:2020dbt; Liu:2021fka; Qi:2023oxv; Liu:2024yib; Abedin:2025yru. The precision of the reconstructed $m_{\rm B}$ or $\mu$ and the width of their confidence intervals depend on hyperparameter choices, such as network architecture and training strategy, which influence the expressive power and generalization capability of ANN. Nonetheless, the ANN‑based reconstruction remains a powerful, model‑independent tool for evaluating the CDDR using BAO and SN data.
We perform Bayesian analysis using the publicly available package Cobaya [^4] https://github.com/CobayaSampler/cobaya. Torrado:2020dgo and assess the convergence of the Markov Chain Monte Carlo (MCMC) Lewis:2002ah; Lewis:2013hha chains using the Gelman-Rubin statistics quantity R1<0.01R-1<0.01 Gelman:1992zz. The MCMC chains are analyzed using GetDist [^5] https://github.com/cmbant/getdist. Lewis:2019xzd.
III Results and discussions
Figure 2: The triangular plots on $\eta_{0}$ and $M_{\rm B}$ for the $P_{1}$ (left panel), $P_{2}$ (middle panel), and $P_{3}$ (right panel) models, using SDSS, DESI, and PantheonPlus data.
Figure 3: The 1D marginalized posterior distributions of the $\eta_{0}$ parameter for the $P_{1}$, $P_{2}$, and $P_{3}$ models of the CDDR with $M_{\rm B}$ fixed, using $M_{\rm B}^{\rm B23}$ (upper panel) and $M_{\rm B}^{\rm D20}$ (lower panel), based on SDSS, DESI, and PantheonPlus data.
In this section, we present the constraints on $\eta_{0}$ and $M_{\rm B}$ for the $P_{1}$, $P_{2}$, and $P_{3}$ parameterizations of $\eta(z)$, using the SDSS, DESI, PantheonPlus, and DESY5 data. The marginalized results with $1\sigma$ errors are summarized in Table 2. The triangular plots, along with the one-dimensional (1D) marginalized distributions of $\eta_{0}$ and $M_{\rm B}$, based on various datasets and parameterizations, are shown in Figs. 23 and 4.
Figure 4: The 1D marginalized posterior distributions of the $\eta_{0}$ parameter for the $P_{1}$, $P_{2}$, and $P_{3}$ models of the CDDR, using SDSS, DESI, and DESY5 data.
In Fig. 2, we show the triangular plot of the constraint results obtained by treating $M_{\rm B}$ as a free parameter, using SDSS, DESI, and PantheonPlus data. The constraints on $\eta_{0}$ are $0.060^{+0.120}_{-0.210}$ ($P_{1}$ model), $0.480^{+0.440}_{-0.990}$ ($P_{2}$ model), and $0.170^{+0.220}_{-0.480}$ ($P_{3}$ model), based on the DESI+PantheonPlus data. This indicates that in all parameterized models, there is no evidence for a violation of the CDDR. This result emphasizes the robustness of our CDDR analysis, and the consistency among the three parameterized models $P_{1}$, $P_{2}$, and $P_{3}$ also suggests that using multiple parameterizations of $\eta(z)$ can provide complementary insights into potential deviations from the CDDR. Similarly, the SDSS+PantheonPlus combination yields consistent results, indicating good agreement between the BAO measurements from SDSS and DESI DR2. The constraints on $M_{\rm B}$ using the DESI+PantheonPlus datasets are $19.560^{+0.420}_{-0.290}$ ($P_{1}$ model), $19.810^{+0.790}_{-0.570}$ ($P_{2}$ model), and $19.650^{+0.570}_{-0.390}$ ($P_{3}$ model). The central values of $M_{\rm B}$ in all parameterized models are slightly lower than the commonly accepted value of approximately 19.3-19.3 mag. This slight shift may result from the inclusion of the possible deviation parameter $\eta_{0}$ in the CDDR analysis.
It is worth noting that there is a strong negative correlation between $\eta_{0}$ and $M_{\rm B}$, as shown in Fig. 2. This correlation indicates that different $M_{\rm B}$ priors will have a significant impact on our detection of potential deviations from the CDDR. Therefore, we further adopt the $M_{\rm B}^{\rm D20}$ and $M_{\rm B}^{\rm B23}$ priors to assess the potential impact of $M_{\rm B}$ on deviations from the CDDR. The 1D marginalized posterior distributions of $\eta_{0}$ for $P_{1}$, $P_{2}$, and $P_{3}$ models are shown in Fig. 3. When adopting the $M_{\rm B}^{\rm D20}$ prior with $M_{\rm B}=-19.230\pm 0.040$ mag, $P_{1}$, $P_{2}$, and $P_{3}$ models yield $\eta_{0}=-0.098\pm 0.052$, $\eta_{0}=-0.220\pm 0.113$, and $\eta_{0}=-0.162\pm 0.068$ using DESI+PantheonPlus data, respectively. All three models consistently show a deviation from CDDR at approximately $2\sigma$ confidence level, especially for the $P_{3}$ model, which has reached $2.4\sigma$. For further discussion on the $\sim 2\sigma$ deviation from the CDDR, please refer to Appendix B. In contrast, if we adopt an $M_{\rm B}^{\rm B23}$ prior with $M_{\rm B}=-19.396\pm 0.016$ mag, the constraint values of $\eta_{0}$ are $0.035\pm 0.056$ ($P_{1}$ model), $0.070\pm 0.124$ ($P_{2}$ model), and $0.043\pm 0.088$ ($P_{3}$ model), suggesting that no deviation from the CDDR is found in all models. This indicates that if a fixed $M_{\rm B}$ prior is adopted, a bias of approximately 0.2 mag of $M_{\rm B}$ can lead to a deviation of $2\sigma$ in CDDR, emphasizing the necessity of analysing the impact of the $M_{\rm B}$ prior on CDDR deviation. Furthermore, we can obtain that the absolute magnitude of the SN needs to be calibrated to approximately 19.4-19.4 mag in order for the CDDR to remain undeviated using the DESI and PantheonPlus data.
Next, we present the 1D marginalized posterior distributions of $\eta_{0}$ for the $P_{1}$, $P_{2}$, and $P_{3}$ models, obtained using the DESI+DESY5 and SDSS+DESY5 data, as shown in Fig. 4. Using the SDSS+DESY5 data, the constraint values of $\eta_{0}$ are $035\pm 0.074, 0.0$$080\pm 0.120$, and $0.059\pm 0.093$ for $P_{1}$, $P_{2}$, and $P_{3}$ models, respectively. Similarly, the constraint values of $\eta_{0}$ are $0.052\pm 0.066$ ($P_{1}$ model), $0.090\pm 0.110$ ($P_{2}$ model), and $0.068\pm 0.085$ ($P_{3}$ model) using DESI+DESY5 data, respectively. We find that these results consistently show no deviation from the CDDR in all parameterized models. Therefore, except for the case of adopting the $M_{\rm B}^{\rm D20}$ prior, we did not detect any deviation from CDDR in other cases, indicating that the current data are well consistent with CDDR.
Finally, it is necessary to compare the constraints on $\eta_{0}$ given by different observations to explore the capabilities of BAO and SN data. The constraints of $\eta_{0}$ using DESI+DESY5 are approximately 36%, 56%, and 47% higher than those obtained using the SDSS eBOSS DR16 BAO dataset combined with the Pantheon SN sample, for $P_{1}$, $P_{2}$, and $P_{3}$ models, respectively Xu:2022zlm. Similarly, the constraints are approximately 49%, 71%, and 72% more restrictive than those obtained from six well-studied strong gravitational lensing systems and PantheonPlus SN data for three models Qi:2024acx. Furthermore, these results are comparable to those inferred from compact radio quasars measurements and Pantheon SN samples Yang:2024icv. Therefore, the latest BAO and SN data provide a more effective and accurate way to test CDDR.
IV Conclusion
The CDDR is a fundamental relationship of modern cosmology, and its validation largely relies on precise distance measurements. In this work, we explore the potential deviation of the CDDR using three phenomenological parameterizations of the function $\eta(z)$, by combining BAO data from SDSS and DESI DR2 with SN data from PantheonPlus and DESY5. We also evaluate the impacts of the SN absolute magnitude on testing the CDDR.
When $M_{\rm B}$ is treated as a free parameter, we find a strong negative correlation between $\eta_{0}$ and $M_{\rm B}$, indicating that the uncertainty in $M_{\rm B}$ directly affects the inferred CDDR deviation. Meanwhile, all three parameterizations show no evidence for a deviation from CDDR using the SDSS, DESI, and PantheonPlus data. The best-fit values of $M_{\rm B}$ are slightly lower than the commonly adopted value of 19.3-19.3 mag, suggesting that the inclusion of a deviation parameter in CDDR can lead to a small shift in $M_{\rm B}$ from its conventional calibration. When fixing the $M_{\rm B}^{\rm D20}$ prior ($19.230\pm 0.040$ mag), we find a deviation from CDDR at the approximately $2\sigma$ level, especially in the $P_{3}$ model, where the deviation reaches the $2.4\sigma$ level. This deviation is due to a calibration tension between $M_{\rm B}^{\mathrm{D20}}$ and the Planck-calibrated $r_{\mathrm{d}}$, a mismatch that is closely connected to the $H_{0}$ tension. In contrast, the $M_{\rm B}^{\rm B23}$ prior ($19.396\pm 0.016$ mag) shows no deviation from CDDR in all models. This implies that a shift of approximately 0.2 mag in $M_{\rm B}$ can lead to a $2\sigma$ level deviation from CDDR. Finally, the joint analysis of DESI, SDSS, and DESY5 data maintains CDDR across all models.
In conclusion, our analysis has conducted a meticulous examination of the CDDR, revealing that the absolute magnitude $M_{\rm B}$ of SN has a significant impact on any deviation from the CDDR. Therefore, it is crucial in future to employ more data to calibrate $M_{\rm B}$ with higher accuracy and precision, thereby determining its value and re‐evaluating the CDDR.
[…]
Acknowledgments
We thank Tian-Yang Sun, Sheng-Han Zhou, Jia-Le Ling, Yi-Min Zhang, and Si-Ren Xiao for their helpful discussions. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12533001, 12575049, 12473001, and 12205039), the China Manned Space Program (Grant No. CMS-CSST-2025-A02), and the National 111 Project (Grant No. B16009).
Data Availability Statement This manuscript has no associated data. [Author’s comment: All data analysed in this work is publicly available and references have been provided in Sect. II.]
Code Availability Statement This manuscript has no associated code/software. [Author’s comment: In this work, we utilized the publicly available software packages Cobaya and GetDist, with relevant references provided in Sect. II]
Appendix A The SN Hubble diagram $D_{\mathrm{L}}(z)$ and the distance duality ratio $\eta(z)$
In this appendix, we present the results of the luminosity distance $D_{\mathrm{L}}(z)$ reconstructed from PantheonPlus (considering two different $M_{\rm B}$ priors) and DESY5 SN data, and the distance duality ratio $\eta(z)$, constructed using $D_{\rm A}$ from BAO and $D_{\rm L}$ from SN, as shown in Fig. 5. The blue data points corresponding to the combination DESI+PantheonPlus $\times$ $M_{\rm B}^{\rm D20}$ and the green data points for SDSS+PantheonPlus $\times$ $M_{\rm B}^{\rm D20}$ exhibit a noticeable deviation from the expected value $\eta(z)=1$, suggesting a potential violation of the CDDR. In contrast, the red data points corresponding to the combination DESI+PantheonPlus $\times$ $M_{\rm B}^{\rm B23}$ and the orange data points for SDSS+PantheonPlus $\times$ $M_{\rm B}^{\rm B23}$ are in better agreement with $\eta(z)=1$. These results underscore the limitations of testing the CDDR using calibration-dependent distance probes, particularly the dependence on the parameter $M_{\rm B}$. Additionally, the purple data points for DESI+DESY5 and the pink data points for SDSS+DESY5 are in close agreement, both being consistent with $\eta(z)=1$.
Appendix B A brief discussion on the $\sim 2\sigma$ deviation from the CDDR
To assess whether the $\sim 2\sigma$ departure from the CDDR when adopting the $M_{\rm B}^{\mathrm{D20}}$ prior is driven by the assumed BAO sound horizon, we repeated the analysis fixing $r_{\rm d}=138.3\ \mathrm{Mpc}$ Wang:2024rxm. The constraint results of $\eta_{0}$ are reported in Table 3. The updated fits show that, for all parametrisations and for both SDSS+PantheonPlus and DESI+PantheonPlus data, the values of $\eta_{0}$ are consistent with zero at less than the $1\sigma$ level; accordingly, the previously reported $\sim 2\sigma$ deviation disappears when a smaller $r_{\mathrm{d}}$ is adopted. This behaviour points to a calibration inconsistency between $M_{\rm B}^{\mathrm{D20}}$ and the Planck-calibrated $r_{\mathrm{d}}$, a mismatch that is also related to the $H_{0}$ tension.
Table 3: Constraint results of $\eta_{0}$ in the $P_{1}$, $P_{2}$, and $P_{3}$ models when $r_{\mathrm{d}}=138.3\ \mathrm{Mpc}$ is adopted Wang:2024rxm.
Model SDSS + PantheonPlus $\times M_{\rm B}^{\rm D20}$ DESI + PantheonPlus $\times M_{\rm B}^{\rm D20}$
$\eta_{0}$ $\eta_{0}$
$P_{1}$ model $043\pm 0.068 0.0$$037\pm 0.052$
$P_{2}$ model $080\pm 0.125 0.0$$080\pm 0.110$
$P_{3}$ model $065\pm 0.089 0.0$$068\pm 0.083$